You are going to invest $25,000, part at 14% and part at 16%. What is the most that can be invested at 14% in order to make at least $3000 interest per year?
Answer by RAY100(1637)

Let x = amt at 14% and y = amt at 16%
.
.14x +.16y = $ 3000
.
but,,,x+y=$ 25,000,,,,,,,or y=25,000 -x,,,,,subst
.
.14x + .16(25,000-x) = 3000
.
.14x + 4,000 - .16x = 3000
.
- .02 x = -1000
.
x= $ 50,000,,,,,y= - $ 25,000
.
Checking,,, 50000(.14) - 25000(.16) = 7000 - 4000 = 3000,,,math correct BUT
.
,
To make $ 3000 from $ 25,000,, need return of 3000/25000= .12 or 12 % ,,,only
.
.
The real question is how much at 14 % = $ 30,000
.
.14 * x = $3000,,,,x= 3000/ .14 = $ 21,428.57
.
.
.
If we invested all at 16%, we would invest,,,y= 3000/.16 = $ 18,750
.